OpenStudy (anonymous):

An 80.0-gram sample of a gas was heated from 25 °C to 225 °C. During this process, 346 J of work was done by the system and its internal energy increased by 6275 J. What is the specific heat of the gas?

OpenStudy (cuanchi):

delta E = q+w q= delta E- w = m x C x deltaT C=delta E- w /(m x deltaT) delta E= 6275 (it is positive the system increase the energy) w= -346J (it is negative the system release work) m=80.0g delta T= Tf-Ti= 225-25= 200 (it is positive the system increase its temperature

OpenStudy (anonymous):

Thank you